OpenStudy (anonymous):

What is square root -112 simplified please help :)

OpenStudy (saifoo.khan):

\[4i \sqrt7\]

OpenStudy (anonymous):

thanks a lot! can you tell me how you got that? or point me to a direction to get help on this?

OpenStudy (saifoo.khan):

Sure.

OpenStudy (anonymous):

-112 is the same as -1*16*7. So take the square roots of each of those

OpenStudy (saifoo.khan):

\[\huge \sqrt{-112} \to\sqrt{-4\times 4\times 7} \to4\sqrt{-1\times7} \to4i \sqrt7\]

OpenStudy (anonymous):

if i used the factor tree method like this..